OpenStudy (anonymous):

OK, I am really lost on this one. AN Airbus 320 plane takes off from Denver International Airport, which is 5200 ft. above sea level, and climbs to 30,000 ft at a constant rate of 1000ft/min. A.) What is the linear equation for the height of the plane in terms of time after take off? B.) Using the above equation find the height of the plane 8 minutes after take off

OpenStudy (radar):

h=5200+1,000t for height after 8 min. h=5200+8,000=13,200 feet

OpenStudy (anonymous):

Thank you radar!